OpenStudy (pokemon23):

A purse containing $3.20 in quarters and dimes has, in all 20 coins. Find the number of each kind of coin. A purse containing $3.20 in quarters and dimes has, in all 20 coins. Find the number of each kind of coin. @Mathematics

OpenStudy (pokemon23):

help me with this question

OpenStudy (lilg132):

8 quarters 12 dimes

OpenStudy (pokemon23):

i need how you got it?

OpenStudy (mertsj):

Let x = the number of quarters Then 20-x = the number of dimes 25x + 10(20-x) = 320 ALL CENTS 25x +200 - 10x = 320 15x = 120 x = 8 quarters 20-x = 12 dimes

OpenStudy (lilg132):

ok x = number of quarters number of dimes = 20 - x 0.25(x) + 0.1(20-x) = 3.20 0.25x + 2 - 0.1x = 3.20 0.25 - 0.1x = 3.20 - 2 0.15x = 1.20 x = 8 quarters 20 - 8 = 12 dimes